Calculate the modulus of elasticity for a composite material consisting of 60 percent
by volume of continuous E-glass fiber and 40 percent epoxy resin for the matrix when
stressed under isostress conditions. The modulus of elasticity of the E glass is 72.4 GPa and
that of the epoxy resin is 3.1GPa.
Answer #1
E = E1昖1 +E2昖2 = 0.6?72.4 +
0.4?3.1 =43.44 +1.24 =44.68 GPa
answered by: Marcie <3s Jeffrey